OpenStudy (anonymous):

You select a number at random from the sample space {1,2,3,4,5}. Find each theoretical probability.

OpenStudy (anonymous):

Probability of selecting 1=1/5 Probability of selecting 2=1/5 Probability of selecting 3=1/5 Probability of selecting 4=1/5 Probability of selecting 5=1/5